OpenStudy (anonymous):

You are working on a new temperature scale that will unify the earth. Let F denote the temperature in degrees Fahrenheit, let X denote your new temperature scale. You want it to be such that if F=0 then X=19 and if F=100 then X=107 . You also want X to be such that if you plot X against F you obtain a straight line. The formula X= mF + b Where M = ????? Where B = ?????

OpenStudy (amistre64):

so they give us 2 "coordinates" to work with, and that is all you really need for a line right?

OpenStudy (amistre64):

0,19 and 100,107 lets find the slope between these points

OpenStudy (amistre64):

107-19 88 ------- = --- = 44/50 = 22/25 100-0 100

OpenStudy (amistre64):

M = 22/25 0 = (22/25)(19) + B 0 = 418/25 + B B = -418/25 simplify as needed :)

OpenStudy (amistre64):

and check as required lol.... i used it backwards

OpenStudy (amistre64):

M = 22/25 19 = (22/25)0 + B 19 = 0 + B B = 19 .... thats better ;)

OpenStudy (anonymous):

Thanks! Wow that worked out well:) that makes sense, thanks!!!